Derik ran 3/4 of a mile each way to and from school Monday through Friday.how far did run that week?


1
Expert's answer
2021-03-01T16:09:53-0500
"\dfrac{3}{4}\cdot2\cdot5=\dfrac{15}{2}(miles)=7\dfrac{1}{2}(miles)"

Derik ran "\dfrac{15}{2}" (or "7\dfrac{1}{2}") miles that week.